# Hoaxpedia: A Unified Wikipedia Hoax Articles Dataset

### v2

@@ Abstract
Hoaxes are a recognised form of disinformation created deliberately, with potential serious implications in the credibility of reference knowledge resources such as Wikipedia. What makes detecting Wikipedia hoaxes hard is that they often are written according to the official style guidelines. In this work, we first provide a systematic analysis of the similarities and discrepancies between legitimate and hoax Wikipedia articles, and introduce Hoaxpedia, a collection of 311 Hoax articles (from existing literature as well as official Wikipedia lists) alongside semantically similar real articles. We report results of binary classification experiments in the task of predicting whether a Wikipedia article is real or hoax, and analyze several settings as well as a range of language models. Our results suggest that detecting deceitful content in Wikipedia based on content alone, despite not having been explored much in the past, is a promising direction. [^1] The Dataset(view-only, access upon request) is available in: https://osf.io/rce8m/?view_only=ed469941644c496fb4a6425297ced1f2. We will publicly release our models and the datasets in Huggingface upon acceptance
@@
Hsuvas Borkakoty1, Luis Espinosa-Anke1,2, 1Cardiff NLP, School of Computer Science and Informatics, Cardiff University, UK 2AMPLYFI, UK {borkakotyh,espinosaankel}@cardiff.ac.uk
@@ 1 Introduction
Wikipedia is, as Hovy et al. (2013) define it, the “largest and most popular collaborative and multilingual resource of world and linguistic knowledge”, and it is acknowledged that its accuracy is on par with or superior than, e.g., the Encyclopedia Britannica Giles (2005). However, as with any other platform, Wikipedia is also the target of online vandalism, and hoaxes, a more obscure, less obvious form of vandalism [^2] https://en.wikipedia.org/wiki/Wikipedia:Do_not_create_hoaxes., constitute a threat to the overall integrity of this collaborative encyclopedia (Kumar et al., 2016; Wong et al., 2021; Wang and McKeown, 2010), precisely because of its “publish first, ask questions later” policy Asthana and Halfaker (2018). Although Wikipedia employs community based New Page Patrol systems to check the credibility of a newly created article, the process is always in backlog [^3] https://en.wikipedia.org/wiki/Wikipedia:New_pages_patrol., making the process overwhelming Schneider et al. (2014).
@@
Hoax articles, created to deliberately spread false information Kumar et al. (2016), harm the credibility of Wikipedia as a knowledge resource, and generate concerns among its users Hu et al. (2007). Since manual inspection of quality is a process typically in backlog Dang and Ignat (2016), the automatic detection of such articles is a desirable feature. However, most works in the literature have centered their efforts in metadata associated with hoax articles, e.g., user activity, appearance features or revision history Zeng et al. (2006); Elebiary and Ciampaglia (2023); Kumar et al. (2016); Wong et al. (2021); Hu et al. (2007); Susuri et al. (2017). For example, Adler et al. (2011) introduced a vandalism detection system using metadata, content and author reputation features, whereas Kumar et al. (2016) provides a comprehensive study of hoax articles and their timeline from discovery to deletion. In this work, the authors define the characteristics of a successful hoax, with a data-driven approach based on studying a dataset of 64 articles (both hoax and real), on top of which they train statistical classifiers. Furthermore, other works have compared network traffic and features of hoax articles to those of other articles published the same day (Elebiary and Ciampaglia, 2023), and conclude that hoax articles attract more attention after creation than cohort articles. Finally, Wong et al. (2021) study various Wikipedia vandalism types and introduce the Wiki-Reliability dataset, which comprises articles based on 41 author-compiled templates. This dataset contains 1,300 articles marked as hoax, which are real articles with false information, a.k.a hoax facts (Kumar et al., 2016).
@@
We part ways from previous works and focus exclusively on the content of hoax articles, and aim to answer the research question “Can hoax articles be distinguished from real articles using NLP techniques by looking exclusively at the article’s content”? We first construct a dataset (HoaxPedia) containing 311 hoax articles and around 30,000 plausible negative examples, i.e., real Wikipedia articles that are semantically similar to hoax articles, to create a set of negatives that cover similar topics to hoax articles (e.g., a newly discovered species). We also explore whether a Wikipedia definition (the first sentence of the article) can provide hints towards its veracity. Our results (reported at different ratios of hoax vs real articles) suggest that, while style and shallow features are certainly not good predictors of hoax vs real Wikipedia articles, LMs are capable of exploiting other more intricate features, and open a promising research direction focused on content-based hoax flagging. Our contributions in this work can be summarised as follows.
@@
  • We systematically contrast a set of proven Wikipedia hoax articles vs. legitimate articles.
@@
  • We propose HoaxPedia, a novel Wikipedia Hoax article dataset with 311 hoax articles and semantically similar legitimate articles collected from Wikipedia.
  • We conduct binary classification experiments on this dataset, using a range of language models to accurately predict whether an article is a hoax or real based solely on its content.
@@ 2 HoaxPedia Construction
HoaxPedia is constructed by unifying five different resources that contain known hoaxes, e.g., from Kumar et al. (2016); Elebiary and Ciampaglia (2023), as well as the official Wikipedia hoaxes list [^4] https://en.wikipedia.org/wiki/Wikipedia:List_of_hoaxes_on_Wikipedia and the Internet Archive. We used Internet Archive to manually retrieve Wikipedia pages that are now deleted from Wikipedia, but were at one point in the past identified as hoaxes. We manually verify each of the article we collect from Wikipedia and Internet Archive as hoax using their accompanied deletion discussion and reasons for citing them as a hoax. In terms of negative examples, while we could have randomly sampled Wikipedia pages, this could have introduced a number of biases in the dataset, e.g., hoax articles contain historical events, personalities or artifacts, and thus we are interested in capturing a similar breadth of topics, entities and sectors in the negative examples so that a classifier cannot rely on these spurious features. We select these negative examples (truthful articles) by ensuring they correspond to authentic content. This is achieved by verifying they do not carry the Db-hoax flag, which Wikipedia’s New Page Patrol policy uses to mark potential hoaxes. Within this set, we extract negative examples as follows. Let $H$ be the set of hoax articles, and $W$ the set of candidate real Wikipedia pages, with $T_{H}=\{t_{H^{1}},\dots,t_{H^{p}}\}$ and $T_{W}=\{t_{W^{1}},\dots,t_{W^{q}}\}$ their corresponding vector representations, and $p$ and $q$ the number of hoax and candidate Wikipedia articles, respectively. Then, for each SBERT(all-MiniLM-L6-v2) ( (Reimers and Gurevych, 2019) title embedding $t_{H^{i}}\in T_{H}$, we retrieve its top $k$ nearest neighbors (NN) from $T_{W}$ via cosine similarity cos. We experiment with different values for $k$, specifically $k\in\{2,10,100\}$:
@@
3 Hoax vs. Real, a Surface-Level Comparison
To maintain the longevity and to avoid detection, hoax articles follow Wikipedia guidelines and article structure. This raises the following question: “how (dis)similar are hoaxes with respect to a hypothetical real counterpart?”. Upon inspection, we found comments in the deletion discussions such as “I wouldn’t have questioned it had I come across it organically” (for the hoax article The Heat is On  [^5] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/The_Heat_Is_On_(TV_series)), or “The story may have a "credible feel” to it, but it lacks any sources”, a comment on article Chu Chi Zui [^6] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/Chu_Chi_Zui. Comments like these highlight that hoaxes are generally well written (following Wikipedia’s guidelines), and so we proceed to quantify their stylistic differences in a comparative analysis that looks at: (1) article text length comparison; (2) sentence and word length comparison; and (3) a readability analysis.
@@
Following the works of Kumar et al. (2016), we conduct a text length distribution analysis with hoax and real articles, and verify they show a similar pattern (as shown in Figure 1), with similar medians for hoax and real articles, specifically 1,057 and 1,777 words, respectively.
Figure 1: Text length distribution for hoax and real articles.
@@
Calculating average sentence and word length for hoaxes and real articles separately can be a valuable proxy for identifying any obvious stylistic or linguistic (e.g., syntactic complexity) patterns. We visualize these in a series of box plots in Figure 2. They clearly show a similar style, with sentence and word length medians at 21.23 and 22.0, and 4.36 and 4.35 for real and hoax articles respectively.
@@
Figure 2: Results of different stylistic analyses on Hoax (red) and Real (blue) articles.
@@
Readability analysis gives a quantifiable measure of the complexities of texts. It can clarify on easement of understanding the text, revealing patterns that can be used to either disguise disinformation like hoaxes or convey clear, factual content. For readability analysis, we use Flesch-Kincaid (FK) Grading system (Flesch, 2007), a metric that indicates the comprehension difficulty when reading a passage in the context of contemporary academic English. This metric gives us an aggregate of the complexity of documents, their sentences and words, measured by the average number of words per sentence and the average number of syllables per word. After obtaining an average for both hoax and real articles, we visualize these averages again in Figure 2, we find a median of 9.4 for real articles and 9.5 for hoax articles, again highlighting the similarities between these articles.

### v3

@@ Abstract
Hoaxes are a recognised form of disinformation created deliberately, with potential serious implications in the credibility of reference knowledge resources such as Wikipedia. What makes detecting Wikipedia hoaxes hard is that they often are written according to the official style guidelines. In this work, we first provide a systematic analysis of similarities and discrepancies between legitimate and hoax Wikipedia articles, and introduce Hoaxpedia, a collection of 311 hoax articles (from existing literature and official Wikipedia lists), together with semantically similar legitimate articles, which together form a binary text classification dataset aimed at fostering research in automated hoax detection. In this paper, We report results after analyzing several language models, hoax-to-legit ratios, and the amount of text classifiers are exposed to (full article vs the article’s definition alone). Our results suggest that detecting deceitful content in Wikipedia based on content alone is hard but feasible, and complement our analysis with a study on the differences in distributions in edit histories, and find that looking at this feature yields better classification results than context.  [^1] The Dataset is available at: https://huggingface.co/datasets/hsuvaskakoty/hoaxpedia and associated codes are available at: https://github.com/hsuvas/hoaxpedia_dataset.git
@@
Hsuvas Borkakoty1, Luis Espinosa-Anke1,2 1Cardiff NLP, School of Computer Science and Informatics, Cardiff University, UK 2AMPLYFI, UK {borkakotyh,espinosaankel}@cardiff.ac.uk
@@ 1 Introduction
Wikipedia is, as Hovy et al. (2013) defines it, the “largest and most popular collaborative and multilingual resource of world and linguistic knowledge”, and it is acknowledged that its accuracy is on par with or superior than, e.g., the Encyclopedia Britannica Giles (2005). However, as with any other platform, Wikipedia is also the target of online vandalism, and hoaxes, a more obscure, less obvious form of vandalism [^2] https://en.wikipedia.org/wiki/Wikipedia:Do_not_create_hoaxes., constitute a significant threat to Wikipedia’s overall integrity (Kumar et al., 2016; Wong et al., 2021; Wang and McKeown, 2010), among others, because of its “publish first, ask questions later” policy Asthana and Halfaker (2018). Although Wikipedia employs community based New Page Patrol systems to check the credibility of a newly created article, the process is always in backlog [^3] https://en.wikipedia.org/wiki/Wikipedia:New_pages_patrol., making it overwhelming Schneider et al. (2014).
Figure 1: Example of Hoax and Legitimate(Real) article
@@
Hoax articles (as shown in Figure 1), are created to deliberately spread false information Kumar et al. (2016), harm the credibility of Wikipedia as a knowledge resource and generate concerns among its users Hu et al. (2007). Since manual inspection of quality is a process typically in backlog Dang and Ignat (2016), the automatic detection of such articles is a desirable feature. However, most works in the literature have centered their efforts in metadata associated with hoax articles, e.g., user activity, appearance features or revision history Zeng et al. (2006); Elebiary and Ciampaglia (2023); Kumar et al. (2016); Wong et al. (2021); Hu et al. (2007); Susuri et al. (2017). For example, Adler et al. (2011) introduced a vandalism detection system using metadata, content and author reputation features, whereas Kumar et al. (2016) provide a comprehensive study of hoax articles and their timeline from discovery to deletion. In their work, the authors define the characteristics of a successful hoax, with a data-driven approach based on studying a dataset of 64 articles (both hoax and legitimate), on top of which they train statistical classifiers. Furthermore, other works have compared network traffic and features of hoax articles to those of other articles published the same day (Elebiary and Ciampaglia, 2023), and conclude that hoax articles attract more attention after creation than cohort (or legitimate) articles. Finally, Wong et al. (2021) study various Wikipedia vandalism types and introduce the Wiki-Reliability dataset, which comprises articles based on 41 author-compiled templates. This dataset contains 1,300 articles marked as hoax, which are legitimate articles with false information, a.k.a hoax facts (Kumar et al., 2016).
@@
In this paper, we propose to incorporate content to the mix of features/signals that are fed to classifiers for the hoax detection task, and aim to answer the research question. We first construct a dataset (HoaxPedia) containing 311 hoax articles and around 30,000 plausible negative examples, i.e., legitimate Wikipedia articles that are semantically similar to hoax articles, to create a set of negatives that cover similar topics to hoax articles (e.g., a newly discovered species). We also explore whether a Wikipedia definition (the first sentence of the article) can provide hints towards its veracity. Our results (reported at different ratios of hoax vs. legitimate articles) suggest that, while style and shallow features are certainly not good the best predictors, language models (LMs) are capable of exploiting other more intricate features (e.g., an article’s revision history) and open a promising research direction focused on content-based hoax flagging. Our contributions in this work can be summarised as follows.
@@
  • We systematically contrast a set of proven Wikipedia hoax articles with legitimate articles.
@@
  • We propose HoaxPedia, a novel Wikipedia Hoax article dataset with 311 hoax articles and 30,000 semantically similar legitimate articles collected from Wikipedia.
  • We conduct binary classification experiments on HoaxPedia, using a range of language models (including LLMs), features and hoax-to-legitimate ratio.
2 Related work
In what follows, we give a brief overview of disinformation detection, the datasets available for the community and the role of Wikipedia in disinformation detection, as our work falls in the intersection between disinformation detection and Wikipedia research.
Disinformation detection and datasets:
Disinformation and misinformation are two types of false information, they differ in that misinformation is inaccurate information created or propagated unknowingly, whereas disinformation is inaccurate information deliberately created to mislead the intended consumer Hernon (1995); Fallis (2014); Kumar et al. (2016); Ireton and Posetti (2018). Nonetheless, both are harmful towards information quality and reliability, thus posing risks toward different aspects of society Su et al. (2020). Alam et al. (2021) surveys disinformation detection from a multi-modal perspective, showing that it can come in different modalities like text, images, audio and video, with text being the most common. Datasets used for disinformation detection can be divided based on the length of input or claim: short sentences (such as tweets or Reddit posts) vs articles (common type being news articles), where most of the datasets follow claim-evidence based format Su et al. (2020). The short sentences or claim based datasets are mostly sourced from social media, such as X (formerly Twitter) Castillo et al. (2011); Derczynski et al. (2017); Zubiaga et al. (2018); López and Madhyastha (2021), Reddit Gorrell et al. (2018); Qu et al. (2022), or fact checking websites like Politifact [^4] https://www.politifact.com/ Wang (2017), Snopes [^5] https://www.snopes.com/ Vo and Lee (2020), or a combination of different fact checking websites Augenstein et al. (2019). These datasets contain claims which need to be verified as well as the response, in the form of True/False labeling backed by evidences, which are also provided in these datasets. Article level datasets, on the other hand, are varied, and focus on state-backed propaganda Heppell et al. (2023), German multi-label disinformation (the GerDISDETECT dataset) Schütz et al. (2024), or narratives at conflict dataset containing news articles Sinelnik and Hovy (2024), which mostly focuses on news article or propaganda based disinformation spreading. The datasets mentioned above are specialized towards topic/trend based or news based disinformation, lacking a generalised aspect of disinformation detection with free text.
Wikipedia in disinformation detection:
Wikipedia, as described by McDowell and Vetter (2020), serves as a source of information validation as backed by its large set of articles contributed by community. This is seen in action for fact verification task datasets such as FEVER Thorne et al. (2018b), TabFactA Chen et al. (2019), FNC-1 (Fake News Challenge-1) dataset Pomerleau and Rao (2017) etc. In these datasets, evidences for claims are collected from Wikipedia articles (eg. FEVER, FNC-1) and tables (eg. TabFactA). However, being a product of community effort, Wikipedia is also prone to vandalism and inaccurate contents McDowell and Vetter (2020), and the community outlines different policies to combat these issues [^6] https://en.wikipedia.org/wiki/Wikipedia:Vandalism. We also find efforts to automatize the process of detecting vandalism contents from Natural Language Processing perspective. Preliminary approaches include using feature based approaches extracted from metadata and editor behaviour to detect vandalism Wu et al. (2010); Javanmardi et al. (2011); Heindorf et al. (2016). Implementation of early warning system based on metadata and editor behavior is found in work of Kumar et al. (2015), where they propose a dataset of page metadata with list of blocked user and unblocked users, and proposes a feature based matrix for classification using auto-encoders. Yuan et al. (2017) proposes an edit history based approach, where they use behaviour of users over time as feature to create the embedding space for multi-source LSTM network Hochreiter and Schmidhuber (1997). However, these approaches do not consider article text as a marker to detect the vandalism. Wikipedia marks hoax articles as form of vandalism Kumar et al. (2016); Thorne et al. (2018a). A comprehensive discussion about different approaches for hoax detection are mentioned in Section 1. However, it is worth noting that existing research on this topic is not sufficient to make meaningful progress towards effective disinformation detection using Wikipedia, most of which can be attributed to unavailability of a dedicated dataset containing hoax articles as well as their metadata. We find a research gap towards developing datasets and tools for hoax detection where article contents are used, with foundation stone laid by works of Kumar et al. (2016) and Wong et al. (2021). Through our work, we aim to address this by introducing Hoaxpedia, where we unify the existing datasets we found based on article text, as well as collect articles from Wikipedia’s list of hoaxes.
@@ 3 HoaxPedia Construction
HoaxPedia is constructed by unifying five different resources that contain known hoaxes, e.g., from Kumar et al. (2016); Elebiary and Ciampaglia (2023), as well as the official Wikipedia hoaxes list [^7] https://en.wikipedia.org/wiki/Wikipedia:List_of_hoaxes_on_Wikipedia and the Internet Archive. We used Internet Archive to manually retrieve Wikipedia pages that are now deleted from Wikipedia, but were at one point in the past identified as hoaxes. We manually verify each of the article we collect from Wikipedia and Internet Archive as hoax using their accompanied deletion discussion and reasons for citing them as a hoax. In terms of negative examples, while we could have randomly sampled Wikipedia pages, this could have introduced a number of biases in the dataset, e.g., hoax articles contain historical events, personalities or artifacts, and thus we are interested in capturing a similar breadth of topics, entities and sectors in the negative examples so that a classifier cannot rely on these spurious features. We select these negative examples by ensuring they correspond to authentic content. This is achieved by verifying they do not carry the Db-hoax flag, which Wikipedia’s New Page Patrol policy uses to mark potential hoaxes. Within this set, we extract negative examples as follows. Let $H$ be the set of hoax articles, and $W$ the set of candidate legitimate Wikipedia pages, with $T_{H}=\{t_{H^{1}},\dots,t_{H^{p}}\}$ and $T_{W}=\{t_{W^{1}},\dots,t_{W^{q}}\}$ their corresponding vector representations, and $p$ and $q$ the number of hoax and candidate Wikipedia articles, respectively. Then, for each SBERT (all-MiniLM-L6-v2) (Reimers and Gurevych, 2019) title embedding $t_{H^{i}}\in T_{H}$, we retrieve its top $k$ nearest neighbors (NN) from $T_{W}$ via cosine similarity cos. We experiment with different values for $k$, specifically $k\in\{2,10,100\}$:
@@
4 Hoax vs. legitimate, a Surface-Level Comparison
To maintain longevity and avoid detection, hoax articles follow Wikipedia guidelines and article structure. This raises the following question: “how (dis)similar are hoaxes with respect to a hypothetical legitimate counterpart?”. Upon inspection, we found comments in the deletion discussions such as “I wouldn’t have questioned it had I come across it organically” (for the hoax article The Heat is On  [^8] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/The_Heat_Is_On_(TV_series)), or “The story may have a "credible feel” to it, but it lacks any sources”, a comment on article Chu Chi Zui [^9] https://en.wikipedia.org/wiki/Wikipedia:Articles_for_deletion/Chu_Chi_Zui. Comments like these highlight that hoaxes are generally well written (following Wikipedia’s guidelines), and so we proceed to quantify their stylistic differences in a comparative analysis that looks at: (1) article text length comparison; (2) sentence and word length comparison; and (3) a readability analysis.
@@
Following the works of Kumar et al. (2016), we conduct a text length distribution analysis with hoax and legitimate articles, and verify they show a similar pattern (as shown in Figure 2), with similar medians for hoax and legitimate articles, specifically 1,057 and 1,777 words, respectively.
Figure 2: Text length distribution for hoax and legitimate articles.
@@
Calculating average sentence and word length for hoaxes and legitimate articles separately can be a valuable proxy for identifying any obvious stylistic or linguistic (e.g., syntactic complexity) patterns. We visualize these in a series of box plots in Figure 3. They clearly show a similar style, with sentence and word length medians at 21.23 and 22.0, and 4.36 and 4.35 for legitimate and hoax articles respectively.
@@
Figure 3: Results of different stylistic analyses on Hoax (red) and legitimate (blue) articles.
@@
Readability analysis gives a quantifiable measure of the complexities of texts. It can clarify on easement of understanding the text, revealing patterns that can be used to either disguise disinformation like hoaxes or convey clear, factual content. For readability analysis, we use Flesch-Kincaid (FK) Grading system (Flesch, 2007), a metric that indicates the comprehension difficulty when reading a passage in the context of contemporary academic English. This metric gives us an aggregate of the complexity of documents, their sentences and words, measured by the average number of words per sentence and the average number of syllables per word. After obtaining an average for both hoax and legitimate articles, we visualize these averages again in Figure 3, we find a median of 9.4 for legitimate articles and 9.5 for hoax articles, again highlighting the similarities between these articles.
Timeline Analysis:
Analysing the revision timelines of hoaxes and legitimate Wikipedia articles can give valuable insights into revision density and pattern that might differentiate the both sets of articles. To that end, we collect the timelines of hoax and legitimate articles (in all three hoax to legitimate ratio mentioned above) and add a timeline dataset to Hoaxpedia. Since some of the hoax articles were deleted from Wikipedia, so we are only able to obtain 164 hoax articles out of 311 in our dataset. With these hoax articles and their semantically similar legitimate counterparts, we construct the timeline dataset of Hoaxpedia. We check the timeline patterns through the use of Bayesian Online Changepoint Detection (BOCPD) Adams and MacKay (2007) with Kernel Density Estimation (KDE) Węglarczyk (2018) to obtain the dense regions for revision histories of hoax and legitimate articles. Figure 4 show a comparative timelines with marked normalized dense region for a hoax and a legitimate article as example. Timelines of hoax article show that they are quite dormant in terms of their number of revisions and their dense regions are mostly around the beginning and end of the article’s timeline, which can be attributed to the new page patrol and deletion discussion of hoaxes respectively. To test this hypothesis further, we divide the revision timelines of hoax and legitimate articles into quartiles (divide the timeline into four equal parts) and count the number of dense regions over total dense regions. The result for each quartile is given in Table 1, which shows that the proportion of dense region happening at the beginning and at the end are much higher for hoax articles than legitimate ones. We also show the normalised histogram distribution of hoax and legitimate (real) articles in Figure 5.
(a) Plot for an example Hoax article
(b) Plot for an example legitimate article
Figure 4: Timeline based dense region plots for hoax and legitimate articles (with start and end lines for each region and normalized densities marked in grey)
Quartile Hoax Real
Q1 0.69 0.75
Q2 0.02 0.17
Q3 0.04 0.22
Q4 0.75 0.42
Table 1: Average distribution of dense regions per quartile (timeline divided into four parts) for hoax and legitimate (real) articles
Figure 5: Histogram of normalized distribution for number of revisions in dense regions for hoax and legitimate (real) article
Since we find difference in timeline patterns for hoax and legitimate articles, we also test the possibility of classifying hoax article using timeline only. We train a Support Vector Machine (SVM) Vapnik (2013) for this purpose, with TF-IDF Sparck Jones (1972) vectorizer fit on revision history binned by months (all revision timestamps converted to bins of MM-YYYY format) to build the timeline vocabulary. The results for Label 1 (Hoax) in F1-score is in Table 2, which shows that the classifier is able to perform well with discerning hoax article from legitimate ones, with an effect of data imbalance clearly showing in model’s performance. This further proves that revision history between hoaxes and legitimate articles are different in its pattern, and can be an important feature in detection of hoaxes. However, we also argue that timeline alone may not be able solve the detection of hoaxes, as it is a statistical feature prone to outliers.
Data Split Precision Recall F1
1H2R 0.86 0.91 0.88
1H10R 0.89 0.78 0.83
1H100R 0.97 0.69 0.80
Table 2: Results of SVM Timeline classifier for Label 1 (Hoax) for all data splits